OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3? y=−3x+29 y=3x+13 y=13x−29 y=−13x−13

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 7}}\quad ,&{\color{blue}{ 8}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

sooo y-7=m(x-8)

OpenStudy (anonymous):

y-8=m(x-7)

OpenStudy (jdoe0001):

yes.. notice that slope = m = -3 so \(\bf y-{\color{blue}{ 8}}={\color{green}{ -3}}(x-{\color{red}{ 7}})\)

OpenStudy (anonymous):

y−8=−3(x−7),so how would i solve that

OpenStudy (jdoe0001):

well, first off, distribute the -3 then add 8 to both sides so \(\bf y-{\color{blue}{ 8}}={\color{green}{ -3}}(x-{\color{red}{ 7}})\implies y-8=-3\cdot x-(-3\cdot 7)\\ \quad \\ y-8=-3x+21 y\cancel{ -8+8 }=-3x+21+8\implies y=-3x+29\)

OpenStudy (anonymous):

oh,ok now i see.Thanks for your help.

OpenStudy (jdoe0001):

hmm looks a bit ... odd anyhow \(\bf y-{\color{blue}{ 8}}={\color{green}{ -3}}(x-{\color{red}{ 7}})\implies y-8=-3\cdot x-(-3\cdot 7) \\ \quad \\ y-8=-3x+21\implies y=-3x+29\) yw